# Law of Large Numbers for continuous $N$-particle ensembles at fixed temperature

### v1

@@ 4.2 Proof of the “if” part of the main theorem
where the last line follows because $|\mu|=\mu_{1}+\dots+\mu_{s}$ and $\mu_{i}$ is the total number of steps of any Lukasiewicz path $\Gamma_{i}$ of length $\mu_{i}$. Finally, when $N\to\infty$, by assumption (a) of Definition 3.2, it follows that Eqn. (31) converges to
$\prod_{i=1}^{s}\left[\sum_{\Gamma_{i}\in\mathbf{L}(\mu_{i})}{\prod_{d\geq 1}{(\kappa_{d})^{\#\text{steps $(1,d-1)$ of }\Gamma}}}\right]=\prod_{i=1}^{s}m_{\mu_{i}},$ (32)
as desired.∎
@@ 5 Necessary conditions for the LLN
In this section, we assume that $\{\mu_{N}\}_{N\geq 1}$ satisfies the fixed-temperature LLN and will prove that the two conditions (a)-(b) from Definition 3.2 are satisfied. The key tool will be the formula of Chapuy–Dołęga [CD22] for the multivariate Bessel function (see Theorem 5.6 below).
@@ 5.1 Preliminaries on constellations
We recall here the definition of a constellation, which is a graph embedded on a closed surface that satisfies certain combinatorial rules. Recall that closed connected surfaces are classified by the Euler characteristic, given by the formula
$\chi=V-E+F,$
where $V$, $E$ and $F$ denote the number of vertices, edges and faces of any embedded graph on the surface. The genus $g$ is the quantity given by
$\chi=2-2g,$
for both orientable and non-orientable surfaces. [^4] There is a convention in the literature to take $\chi=2-k$ for non-orientable surfaces, but it is different from ours.
From hereinafter, all closed surfaces are assumed to be connected. It is well known that, up to homeomorphism, each orientable closed surface is uniquely determined by its genus $g\in\mathbb{Z}_{\geq 0}$, while each non-orientable one is uniquely determined by $g\in\frac{1}{2}\mathbb{Z}_{\geq 1}$.
@@
  • Each vertex of color 0 is only connected to vertices of color 1.
  • Each vertex of color $k$ is only connected to vertices of color k1k-1.
  • For $i=1,2,\dots,k-1$, each corner at a vertex of color $i$ separates vertices of colors i1i-1 and $i+1$.
The size of $\mathbf{M}$, denoted by $|\mathbf{M}|$, is the number of corners of color 0 in $\mathbf{M}$. We say that $\mathbf{M}$ is connected if the underlying graph is connected, and $\mathbf{M}$ is orientable/non-orientable if its underlying surface is orientable/non-orientable, respectively. The genus of $\mathbf{M}$ is the genus of the surface it embeds in.
Note that $k\cdot|\mathbf{M}|$ is the number of edges of $\mathbf{M}$.
@@
By definition, in any $k$-constellation $\mathbf{M}$, each vertex of color $k$ is connected only to vertices of color k1k-1. Thus, we can add a new vertex of color $k+1$ for each corner of color $k$ and join it with an edge to the corresponding vertex of color $k$ to obtain a $(k+1)$-constellation of the same size. This map, from $k$-constellations of size $d$, to $(k+1)$-constellations of the same size, preserves rooted and connected constellations. Let $C_{d,\bullet}^{(k)}$ be the set of rooted, connected $k$-constellations of size $d$; the previous discussion thus describes inclusion maps
$i_{k}\colon C_{d,\bullet}^{(k)}\hookrightarrow C_{d,\bullet}^{(k+1)},$
for all $k\geq 1$. A rooted, connected, infinite constellation of size $d$ is any sequence $(\mathbf{M}^{(1)},\mathbf{M}^{(2)},\cdots)$, such that $\mathbf{M}^{(k)}\in C_{d,\bullet}^{(k)}$ and $i_{k}\big(\mathbf{M}^{(k)}\big)=\mathbf{M}^{(k+1)}$, for all $k\geq 1$. It is convenient to think of an infinite constellation simply as the embedding of an infinite graph on a surface, obtained from the construction above (thus, with very “long leaves”). We denote the set of rooted, connected, infinite constellations by $C_{d,\bullet}^{(\infty)}$.
@@
Let $\mathbf{M}$ be an infinite constellation. We define the following notations.
  • For any $f\in F(\mathbf{M})$, we denote by $\deg(f)$ the number of corners of color 0 in $f$.
  • For any $v\in V(\mathbf{M})$, we denote by $\deg(v)$ the number of edges adjacent to vv. [^5] Note that $\deg(v)$ is not necessarily the number of vertices adjacent to $v$, because multiple edges are allowed.
  • We denote by $V_{0}(\mathbf{M})$ the set of vertices of $\mathbf{M}$ of color 0.
  • We denote by $\boldsymbol{\mu_{1}}(\mathbf{M})$ the partition with parts $\deg(f)$, $f\in F(M)$.
  • We denote by $\boldsymbol{\mu_{2}}(\mathbf{M})$ the partition with parts $\deg(v)$, $v\in V_{0}(M)$.
  • For any $i\geq 0$, denote by $v_{i}(\mathbf{M})$ the number of vertices of $\mathbf{M}$ of color $i$.
By definition, note that $|\boldsymbol{\mu_{1}}(\mathbf{M})|=|\boldsymbol{\mu_{2}}(\mathbf{M})|=|\mathbf{M}|$, where $d=|\mathbf{M}|$. Furthermore, denote
$\eta(\mathbf{M}):=(d-v_{i}(\mathbf{M}))_{i\geq 1}.$
We say that $\mathbf{M}$ is normal, if $\eta_{1}(\mathbf{M})\geq\eta_{2}(\mathbf{M})\geq\dots$, so that $\eta(\mathbf{M})$ is a partition.
@@
[CD22, Eqn. (5)] For any infinite constellation $\mathbf{M}$ of genus $g\in\{0,\frac{1}{2},1,\frac{3}{2},\dots\}$, we have
@@
Take a large enough $k$ so that $\eta_{i}(\mathbf{M})=0$, for all $i\geq k$; Moreover, let $d:=|\boldsymbol{\mu_{1}}(\mathbf{M})|=|\boldsymbol{\mu_{2}}(\mathbf{M})|$. Then identify $\mathbf{M}$ as a k-constellation. It follows that the number of edges of $\mathbf{M}$ is $k\cdot d$, the number of vertices of $\mathbf{M}$ is $\sum_{i=0}^{k}{v_{i}(\mathbf{M})}=\sum_{i=1}^{k}{v_{i}(\mathbf{M})}+\ell(\boldsymbol{\mu_{2}}(\mathbf{M}))$, and the number of faces of $\mathbf{M}$ is $\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))$. Since
EVF=2g2,E-V-F=2g-2, (34)
the result follows. ∎
5.2 Chapuy–Dołęga’s expansion of the multivariate Bessel function
Consider the sequences of variables $\mathbf{p}=(p_{1},p_{2},\dots)$, $\mathbf{q}=(q_{1},q_{2},\dots)$, and $(g_{1},g_{2},\dots)$. Let $J_{\lambda}(\mathbf{p};\theta)$ be the Jack symmetric function defined by [Mac98, Ch. VI.10, Eqn. (10.13)], where $\mathbf{p}$ is viewed as the sequence of power sum symmetric functions. Note that $J_{\lambda}(\mathbf{p};\theta)$ is then simply a polynomial of degree $|\lambda|$, if each $p_{k}$ is regarded as a variable of degree $k$. Similarly, we can define $J_{\lambda}(\mathbf{q};\theta)$. Further, consider the generating series $G(z)=1+\sum_{k\geq 1}g_{k}z^{k}$. Following [CD22], define the formal power series666[CD22] uses the parameter $\alpha=\theta^{-1}$ for the Jack parameter instead.
$\hat{F}^{G}(t,\mathbf{p},\mathbf{q},g_{1},g_{2},...,):=\sum_{d\geq 0}t^{d}\sum_{|\lambda|=d}J_{\lambda}(\mathbf{p};\theta)J_{\lambda}(\mathbf{q};\theta)\\ \prod_{(i,j)\in\lambda}\frac{\big(\alpha(\lambda_{i}-j)+(\lambda_{j}^{\prime}-i)+1\big)}{\big(\alpha(\lambda_{i}-j)+(\lambda_{j}^{\prime}-i)+\alpha\big)}\prod_{(i,j)\in\lambda}G(c_{\alpha}(i,j)),$ (35)
where $\alpha=\theta^{-1}$; $\lambda^{\prime}$ is the conjugate partition of $\lambda$; and
[…]
@@ 5.2 Chapuy–Dołęga’s expansion of the multivariate Bessel function
[…]
$\nu_{\rho}\colon\bigcup_{d\geq 1}C_{d,\bullet}^{(\infty)}\to\mathbb{Z}_{\geq 0},$
such that $\nu_{\rho}(\mathbf{M},c)=0$ if and only if $\mathbf{M}$ is orientable ([CD22, Definition-Lemma 3.8]), and such that we have the following equality of formal power series:
$\ln\Big(\hat{F}^{G}\Big)=\theta\cdot\sum_{d\geq 1}\frac{t^{d}}{d}\sum_{(\mathbf{M},\mathbf{c})\in C^{(\infty)}_{d,\bullet},\ \mathbf{M}\ \text{normal}}\big(\theta^{-1}-1\big)^{\nu_{\rho}(\mathbf{M},\mathbf{c})}p_{\boldsymbol{\mu_{1}}(\mathbf{M})}q_{\boldsymbol{\mu_{2}}(\mathbf{M})}\cdot f_{\eta(\mathbf{M})}\left(g_{1},g_{2},\dots\right),$ (36)
where $f_{\eta}(e_{1},e_{2},\dots)$ is the polynomial that expresses the monomial symmetric function $m_{\eta}$ in terms of the elementary symmetric polynomials $e_{1},e_{2},\dots$, and $(\mathbf{M},\mathbf{c})$ in the inner sum ranges over all infinite, rooted, connected constellations of size $d$ that are normal.
@@
Fix $n\in\mathbb{N}$, $\theta>0$, and set $\alpha=\theta^{-1}$. Let $N\in\mathbb{N}$ be large enough so that $N>\max(\alpha,1)\cdot n$. Then, as an analytic function of $(\vec{x},a)\in\mathbb{C}^{N}\times\mathbb{C}^{N}$, we have the expansion
$\begin{split}\ln\Big(B_{a}(\vec{x};\theta)\Big)&\\ =\sum_{d=1}^{n}\frac{1}{d}&\ \sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet},\ \mathbf{M}\ \text{normal},\ g(\mathbf{M})=g}\alpha^{d-1}(\alpha-1)^{\nu_{\rho}(\mathbf{M},c)}\prod_{f\in F(\mathbf{M})}\frac{p_{deg(f)}(\vec{x})}{N}\\ &\times\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\cdot N^{2-2g}\cdot f_{\eta(\mathbf{M})}\Big(\big\{(-1)^{k}\big\}_{k\geq 1}\Big)+O\big(\|x\|^{n+1}\big),\end{split}$ (37)
where $f_{\eta}$ and $(\mathbf{M},\mathbf{c})$ are defined in the same way as in Theorem 5.5, and $g(\mathbf{M})$ denotes the genus of $\mathbf{M}$. The sum in the right hand side converges absolutely.
@@
Recall that the multivariate Bessel functions $B_{a}(\vec{x};\theta)$ are analytic on $\mathbb{C}^{N}\times\mathbb{C}^{N}$ and can be expanded in terms of Jack polynomials as follows (see [OO97, Sec. 4]):
$B_{a}(\vec{x};\theta)=\sum_{\ell(\lambda)\leq N}\frac{\prod_{(i,j)\in\lambda}\big(\alpha(\lambda_{i}-j)+(\lambda_{j}^{\prime}-i)+1\big)}{\prod_{(i,j)\in\lambda}\big(\alpha(\lambda_{i}-j)+(\lambda_{j}^{\prime}-i)+\alpha\big)\big(N\theta+j-1-\theta(i-1)\big)}J_{\lambda}(\vec{x};\theta)J_{\lambda}(a;\theta).$ (38)
Then we specialize $\hat{F}^{G}$ by taking the sequence $g_{k}=(-N)^{-k}$, or equivalently $G(z)=(1+\frac{z}{N})^{-1}$, and let $t=\frac{\alpha}{N}$. Furthermore, for $\vec{x}=(x_{1},...,x_{N})\in\mathbb{C}^{N}$, $a=(a_{1},...,a_{N})\in\mathbb{C}^{N}$, we specify $p_{k}$ and $q_{k}$ as $p_{k}(\vec{x})=\sum_{i=1}^{N}{x_{i}^{k}}$, $q_{k}(a)=\sum_{i=1}^{N}a_{i}^{k}$. By comparing Eqns. (38) and (35), we have
$\begin{split}\hat{F}^{(1+\frac{z}{N})^{-1}}\left(\frac{\alpha}{N},\mathbf{p}(\vec{x}),\mathbf{q}(a),\left\{\left(-N\right)^{-k}\right\}_{k\geq 1}\right)=B_{a}(\vec{x};\theta),\end{split}$ (39)
@@
as an equality of analytic functions, for any $N\in\mathbb{N}$.
By Theorem 5.5, we have the formal expansion:
@@
$\begin{split}\ln\left(B_{a}(\vec{x};\theta)\right)=&\sum_{d=1}^{n}\frac{1}{d}\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet},\ \mathbf{M}\ \text{normal}}\alpha^{d-1}(\alpha-1)^{\nu_{\rho}(\mathbf{M},c)}\prod_{f\in F(\mathbf{M})}\frac{p_{deg(f)}(\vec{x})}{N}\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\\ &\times N^{\ell\big(\boldsymbol{\mu_{1}}(\mathbf{M})\big)+\ell\big(\boldsymbol{\mu_{2}}(\mathbf{M})\big)-\big|\eta(\mathbf{M})\big|}\cdot f_{\eta(\mathbf{M})}\Big(\big\{(-1)^{-k}\big\}_{k\geq 1}\Big)+O\big(\|x\|^{n+1}\big).\end{split}$ (40)
@@
By Proposition 5.4, the exponent of $N$ in the second line of (40) is equal to $\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))+\ell(\boldsymbol{\mu_{2}}(\mathbf{M}))-|\eta(\mathbf{M})|=2-2g$. With given data $\eta(\mathbf{M})$, $\mathbf{M}$ can be identified as a $k$-constellation with $k=\ell(\eta(\mathbf{M}))+1\leq|\eta(\mathbf{M})|+1$. In addition, by [BD22, Theorem 3.1], the total number of $k$-constellation of size $d$ is equal to the total number of $(k+2)$-tuples of matchings(i.e, pairings) of $\{1,\hat{1},2,\hat{2},...,d,\hat{d}\}$, and therefore is finite. We conclude that Eqn. (40) can be rewritten as Eqn. (37), such that for each fixed $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$, $\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet},\ \mathbf{M}\ \text{normal},\ g(\mathbf{M})=g}$ is a finite sum.
@@
It remains to check that when $N>\max(\theta^{-1},1)\cdot n$, the sum over $g$ in (37) is absolute convergent. Using Eqn. (38) and the fact that $B_{a}(0^{N};\theta)=1$, one can expand $\ln\left(B_{a}(\vec{x};\theta)\right)$ in terms of products of the form $\prod_{j=1}^{m}J_{\lambda_{j}}(a;\theta)J_{\lambda_{j}}(\vec{x};\theta)$, for some partitions $\lambda_{1},\dots,\lambda_{m}$, and subsequently, in terms of products of the form $p_{\mu_{1}}(\vec{x})p_{\mu_{2}}(a)$, where $|\mu_{1}|=|\mu_{2}|$. The coefficients in this expansion are rational functions of $N$ that admit a Taylor expansion at $N=\infty$, if $N>\max(\theta^{-1},1)\cdot n$. By Eqn. (39), such expansion agrees with
@@
$\frac{1}{d}\sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}\sum_{(\mathbf{M},c)\in C^{(\infty)}_{d,\bullet},\ \mathbf{M}\ \text{normal},\ g(\mathbf{M})=g,\ \boldsymbol{\mu_{1}}(\mathbf{M})=\mu_{1},\ \boldsymbol{\mu_{2}}(\mathbf{M})=\mu_{2}}\alpha^{d-1}(\alpha-1)^{\nu_{\rho}(\mathbf{M},c)}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\times N^{2-2g-\ell(\mu_{1})-\ell(\mu_{2})-d}\cdot f_{\eta(\mathbf{M})}\Big(\big\{(-1)^{-k}\big\}_{k\geq 1}\Big).$

### v2

@@ 4.2 Proof of the “if” part of the main theorem
Finally, note that for any Łukasiewicz path $\Gamma$, we have
$\text{steps $(1,-1)$ of }\Gamma\,=\,\sum_{d\geq 1}{(d-1)\cdot\#\text{steps $(1,d-1)$ of }\Gamma}.$
Indeed, this equality simply says that the downward $y$-displacement of $\Gamma$ is equal to the upward $y$-displacement of $\Gamma$, but this is true because $\Gamma$ begins at $(0,0)$ and ends at $(\ell,0)$, both points belonging to the $x$-axis. This equality implies that the right hand sides of (31) and (18) coincide. Hence, together with (30)-(31), we deduce the desired equation (29), thus concluding the proof of the “if part” of the main theorem.
@@ 5 Necessary conditions for the LLN
In this section, we assume that $\{\mu_{N}\}_{N\geq 1}$ satisfies the fixed-temperature LLN and prove that the conditions of Definition 3.2 (namely, $\theta$-LLN-appropriateness) are satisfied.
@@ 5.1 Preliminaries on constellations
All our closed surfaces will be connected. Closed surfaces are classified by their genus, which satisfies
22g=VE+F,2-2g=V-E+F,
where $V$, $E$ and $F$ denote the number of vertices, edges and faces of any embedded graph on the surface; moreover, $g$ is the genus of the surface. Up to homeomorphism, the genera of orientable closed surfaces are nonnegative integers, while the genera of non-orientable surfaces are nonnegative half-integers.
@@
$\bullet$ Each vertex of color 0 is only connected to vertices of color 1.
$\bullet$ Each vertex of color $k$ is only connected to vertices of color k1k-1.
$\bullet$ For $1\leq i\leq k-1$, each corner at a vertex of color $i$ separates vertices of colors i1i-1 and $i+1$.
Finally, we will color all corners, too, with the same colors as the ones of the corresponding vertices.
The size of $\mathbf{M}$, denoted by $|\mathbf{M}|$, is the number of corners of color 0 in $\mathbf{M}$. Note that the number of edges of any $k$-constellation $\mathbf{M}$ is $k\cdot|\mathbf{M}|$.
We say that $\mathbf{M}$ is connected if the underlying graph is connected, and $\mathbf{M}$ is orientable/non-orientable if its underlying surface is orientable/non-orientable. The genus of $\mathbf{M}$, denoted $g(\mathbf{M})$, is the genus of the underlying surface.
@@
By definition, in any $k$-constellation $\mathbf{M}$, each vertex of color $k$ is connected only to vertices of color k1k-1. Thus, we can add a new vertex of color $k+1$ for each corner of color $k$ and join it with an edge to the corresponding vertex of color $k$ to obtain a $(k+1)$-constellation of the same size. This size-preserving map from $k$-constellations to $(k+1)$-constellations preserves rooted, connected constellations. Let $C_{d,\bullet}^{(k)}$ be the set of rooted, connected $k$-constellations of size $d$; the previous discussion describes inclusion maps
$i_{k}\colon C_{d,\bullet}^{(k)}\hookrightarrow C_{d,\bullet}^{(k+1)},\quad k\geq 1.$
A rooted, connected, infinite constellation of size $d$ is any sequence $(\mathbf{M}^{(1)},\mathbf{M}^{(2)},\cdots)$ such that $\mathbf{M}^{(k)}\in C_{d,\bullet}^{(k)}$ and $i_{k}\big(\mathbf{M}^{(k)}\big)=\mathbf{M}^{(k+1)}$, for all $k\geq 1$. It is convenient to think of an infinite constellation as the embedding of an infinite graph on a surface, obtained from the construction above. We denote the set of rooted, connected, infinite constellations by $C_{d,\bullet}^{(\infty)}$.
@@
For $\mathbf{M}\in C_{d,\bullet}^{(\infty)}$, let $F(\mathbf{M}),V(\mathbf{M})$ be the sets of faces and vertices of $\mathbf{M}$. Moreover:
$\bullet$ For any $f\in F(\mathbf{M})$, denote by $\deg(f)$ the number of corners of color 0 in $f$. Also, denote by $\boldsymbol{\mu_{1}}(\mathbf{M})$ the partition with parts $\deg(f)$, $f\in F(\mathbf{M})$.
$\bullet$ Denote by $V_{0}(\mathbf{M})$ the set of vertices of $\mathbf{M}$ of color 0, and $v_{0}(\mathbf{M}):=|V_{0}(\mathbf{M})|$.
$\bullet$ For any $v\in V(\mathbf{M})$, denote by $\deg(v)$ the number of edges adjacent to vv. [^2] Note that $\deg(v)$ is not necessarily the number of vertices adjacent to $v$, because multiple edges are allowed. Also, denote by $\boldsymbol{\mu_{2}}(\mathbf{M})$ the partition with parts $\deg(v)$, $v\in V_{0}(\mathbf{M})$.
$\bullet$ For any $i\geq 1$, denote by $v_{i}(\mathbf{M})$ the number of vertices of $\mathbf{M}$ of color $i$. Consider the sequence
$\eta(\mathbf{M}):=\big(|\mathbf{M}|-v_{i}(\mathbf{M})\big)_{i\geq 1}.$
We say that $\mathbf{M}$ is normal if $v_{1}(\mathbf{M})\leq v_{2}(\mathbf{M})\leq\cdots$. Since $v_{i}(\mathbf{M})\leq|\mathbf{M}|$, for all $i$, normality is equivalent to the fact that $\eta(\mathbf{M})$ is a partition.
By the definitions, note that $|\boldsymbol{\mu_{1}}(\mathbf{M})|=|\boldsymbol{\mu_{2}}(\mathbf{M})|=|\mathbf{M}|$.
@@
[CD22, Eqn. (5)] For any $\mathbf{M}\in C_{d,\bullet}^{(\infty)}$ of genus $g$, we have
@@
Take a large enough $k$, so that $\eta_{i}(\mathbf{M})=0$, for all $i\geq k$. Then we can regard $\mathbf{M}$ as a k-constellation. It follows that the number of edges of $\mathbf{M}$ is $k\cdot|\mathbf{M}|$, the number of vertices of $\mathbf{M}$ is $\sum_{i=0}^{k}{v_{i}(\mathbf{M})}=\sum_{i=1}^{k}{v_{i}(\mathbf{M})}+\ell(\boldsymbol{\mu_{2}}(\mathbf{M}))$, and the number of faces of $\mathbf{M}$ is $\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))$. Since EVF=2g2E-V-F=2g-2, the result follows. ∎
5.2 An analytic expansion of the multivariate Bessel function
Consider the sequences of variables $\mathbf{p}=(p_{1},p_{2},\dots)$, $\mathbf{q}=(q_{1},q_{2},\dots)$, and $(g_{1},g_{2},\dots)$. Let $J_{\lambda}(\mathbf{p};\theta)$ be the Jack symmetric function defined by [Mac98, Ch. VI.10, Eqn. (10.13)], where $\mathbf{p}$ is viewed as the sequence of power sum symmetric functions. Then $J_{\lambda}(\mathbf{p};\theta)$ is a polynomial of degree $|\lambda|$, if each $p_{k}$ is regarded as a variable of degree $k$. Define $J_{\lambda}(\mathbf{q};\theta)$ analogously. Further, consider the generating series $G(z)=1+\sum_{k\geq 1}g_{k}z^{k}$. Following [CD22], define the formal power series
$\hat{F}^{G}(t,\mathbf{p},\mathbf{q},g_{1},g_{2},\dots;\theta):=\sum_{d\geq 0}t^{d}\sum_{|\lambda|=d}J_{\lambda}(\mathbf{p};\theta)J_{\lambda}(\mathbf{q};\theta)\\ \prod_{(i,j)\in\lambda}\frac{\big((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta\big)}{\big((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+1\big)}\prod_{(i,j)\in\lambda}G(c_{\theta}(i,j)),$ (33)
where $\lambda^{\prime}_{j}$ denotes the number of boxes of the $j$-th column of $\lambda$ and $c_{\theta}(i,j):=\theta^{-1}(j-1)-(i-1)$.
@@ 5.2 An analytic expansion of the multivariate Bessel function
[CD22, Theorem 6.2] There exists a map [^3] The map $\nu$ measures the “non-orientability” of a connected rooted constellation $\mathbf{M}$: see [CD22, Section 3].
$\nu\colon\bigcup_{d\geq 1}C_{d,\bullet}^{(\infty)}\to\mathbb{Z}_{\geq 0},$
such that $\nu(\mathbf{M})=0$ if and only if $\mathbf{M}$ is orientable ([CD22, Definition-Lemma 3.8]), and such that we have the following equality of formal power series:
$\ln\Big(\hat{F}^{G}\Big)=\theta\cdot\sum_{d\geq 1}\frac{t^{d}}{d}\sum_{\mathbf{M}\in C^{(\infty)}_{d,\bullet}:\ \mathbf{M}\text{ normal}}\big(\theta^{-1}-1\big)^{\nu(\mathbf{M})}p_{\boldsymbol{\mu_{1}}(\mathbf{M})}q_{\boldsymbol{\mu_{2}}(\mathbf{M})}\cdot f_{\eta(\mathbf{M})}\left(g_{1},g_{2},\dots\right),$ (34)
where $f_{\eta}(e_{1},e_{2},\dots)$ is the polynomial that expresses the monomial symmetric function $m_{\eta}$ in terms of the elementary symmetric polynomials $e_{1},e_{2},\dots$, and $\mathbf{M}$ in the inner sum ranges over all rooted, connected, normal, infinite constellations of size $d$.
@@
Fix $n\in\mathbb{N}$, $\theta>0$, and set $\alpha=\theta^{-1}$. Let $N\in\mathbb{N}$ be large enough so that $N>\max(\alpha,1)\cdot n$. Then, as an analytic function of $(a,\vec{x})\in\mathbb{C}^{N}\times\mathbb{C}^{N}$, we have the expansion
$\ln\Big(B_{a}(\vec{x};\theta)\Big)=\sum_{d=1}^{n}\frac{\alpha^{d-1}}{d}\sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}N^{2-2g}\sum_{\begin{subarray}{c}\mathbf{M}\in C^{(\infty)}_{d,\bullet}\\ \mathbf{M}\text{ normal},\ g(\mathbf{M})=g\end{subarray}}(\alpha-1)^{\nu(\mathbf{M})}\prod_{f\in F(\mathbf{M})}\frac{p_{deg(f)}(\vec{x})}{N}\\ \times\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\cdot f_{\eta(\mathbf{M})}\Big(\big\{(-1)^{k}\big\}_{k\geq 1}\Big)+O\big(\|x\|^{n+1}\big),$ (35)
where $f_{\eta}$ and $\mathbf{M}$ are defined in the same way as in Theorem 5.5. The sum over $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ in the right hand side converges absolutely.
@@
The multivariate Bessel functions $B_{a}(\vec{x};\theta)$ are analytic functions of $(a,\vec{x})\in\mathbb{C}^{N}\times\mathbb{C}^{N}$ and can be expanded in terms of Jack polynomials as follows (see [OO97, Sec. 4]):
$B_{a}(\vec{x};\theta)=\sum_{\ell(\lambda)\leq N}\frac{\prod_{(i,j)\in\lambda}\big((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+\theta\big)}{\prod_{(i,j)\in\lambda}\big((\lambda_{i}-j)+\theta(\lambda_{j}^{\prime}-i)+1\big)\big(N\theta+j-1-\theta(i-1)\big)}\,J_{\lambda}(\vec{x};\theta)J_{\lambda}(a;\theta).$ (36)
On the other hand, consider the power series $\hat{F}^{G}$ obtained from specializing to the sequence $g_{k}=(-N)^{-k}$, or equivalently, from $G(z)=(1+\frac{z}{N})^{-1}$, and setting $t=(N\theta)^{-1}$. Furthermore, for $\vec{x}=(x_{1},\dots,x_{N})\in\mathbb{C}^{N}$, $a=(a_{1},\dots,a_{N})\in\mathbb{C}^{N}$, we specify $p_{k}$ and $q_{k}$ as $p_{k}(\vec{x})=\sum_{i=1}^{N}{x_{i}^{k}}$, $q_{k}(a)=\sum_{i=1}^{N}{a_{i}^{k}}$. By comparing equations (33) and (36), we have
$\begin{split}\hat{F}^{(1+\frac{z}{N})^{-1}}\left(\frac{1}{N\theta},\mathbf{p}(\vec{x}),\mathbf{q}(a),\left\{\left(-N\right)^{-k}\right\}_{k\geq 1};\theta\right)=B_{a}(\vec{x};\theta),\end{split}$ (37)
@@
as an equality of analytic functions, for any $N\in\mathbb{N}$. By Theorem 5.5, we have the formal expansion:
@@
$\ln\big(B_{a}(\vec{x};\theta)\big)=\sum_{d=1}^{n}\frac{\alpha^{d-1}}{d}\sum_{\mathbf{M}\in C^{(\infty)}_{d,\bullet}\colon\mathbf{M}\text{ normal}}(\alpha-1)^{\nu(\mathbf{M})}\prod_{f\in F(\mathbf{M})}\frac{p_{deg(f)}(\vec{x})}{N}\prod_{v\in V_{0}(\mathbf{M})}\frac{p_{deg(v)}(a/N)}{N}\\ \times N^{\ell\big(\boldsymbol{\mu_{1}}(\mathbf{M})\big)+\ell\big(\boldsymbol{\mu_{2}}(\mathbf{M})\big)-\big|\eta(\mathbf{M})\big|}\cdot f_{\eta(\mathbf{M})}\Big(\big\{(-1)^{-k}\big\}_{k\geq 1}\Big)+O\big(\|x\|^{n+1}\big).$ (38)
@@
By Proposition 5.4, the exponent of $N$ in the second line of (38) is equal to $2-2g(\mathbf{M})$. For a fixed $g(\mathbf{M})=g$, we have that $\ell(\eta(\mathbf{M}))\leq|\eta(\mathbf{M})|=\ell(\boldsymbol{\mu_{1}}(\mathbf{M}))+\ell(\boldsymbol{\mu_{2}}(\mathbf{M}))+2g-2<|\boldsymbol{\mu_{1}}(\mathbf{M})|+|\boldsymbol{\mu_{2}}(\mathbf{M})|+2g=2|\mathbf{M}|+2g=2d+2g\leq 2n+2g$. As a result, if we set $k:=2n+2g$, then $\mathbf{M}$ can be regarded as a $k$-constellation. In addition, by [BD22, Theorem 3.1], the total number of $k$-constellation of size $d$ is finite. We conclude that Eqn. (38) can be rewritten as Eqn. (35), such that for each fixed $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$, the sum $\sum_{\mathbf{M}\in C^{(\infty)}_{d,\bullet}\colon\mathbf{M}\text{ normal},\ g(\mathbf{M})=g}$ is finite.
@@
It remains to check that when $N>\max(\alpha,1)\cdot n$, the sum over $g\in\frac{1}{2}\mathbb{Z}_{\geq 0}$ in (35) converges absolutely. Using Eqn. (36) and the fact that $B_{a}(0^{N};\theta)=1$, one can expand $\ln\left(B_{a}(\vec{x};\theta)\right)$ in terms of products of the form $\prod_{j=1}^{m}J_{\lambda_{j}}(a;\theta)J_{\lambda_{j}}(\vec{x};\theta)$, for some partitions $\lambda_{1},\dots,\lambda_{m}$, and subsequently, in terms of products of the form $p_{\mu_{1}}(\vec{x})p_{\mu_{2}}(a)$, where $|\mu_{1}|=|\mu_{2}|$. The coefficients in this expansion are rational functions of $N$ that admit a Taylor expansion at $N=\infty$, if $N>\max(\alpha,1)\cdot n$. By Eqn. (37), such expansion agrees with
@@
$\frac{\alpha^{d-1}}{d}\sum_{g\in\frac{1}{2}\mathbb{Z}_{\geq 0}}\sum_{\begin{subarray}{c}\mathbf{M}\in C^{(\infty)}_{d,\bullet}\colon\mathbf{M}\text{ normal},\\ g(\mathbf{M})=g,\ \boldsymbol{\mu_{1}}(\mathbf{M})=\mu_{1},\ \boldsymbol{\mu_{2}}(\mathbf{M})=\mu_{2}\end{subarray}}(\alpha-1)^{\nu(\mathbf{M})}N^{2-2g-\ell(\mu_{1})-\ell(\mu_{2})-d}\cdot f_{\eta(\mathbf{M})}\Big(\big\{(-1)^{-k}\big\}_{k\geq 1}\Big).$